What is the probability of drawing a red face card from a deck of cards?
There are 6 red face cards in a standard deck of 52 cards; the Jack, Queen, and King of Hearts and Diamonds. The probability, then, of drawing a red face card from a standard deck of 52 cards is 6 in 52, or 3 in 26, or about 0.1154.
5/16
6 Boys
10 Girls
Half the girls wear glasses. Half of 10 is 5.
5 Girls wear glasses.
Probability=outcome wanted/total
5 girls wear glasses/16 kids in class
or 5/16
Message me with any questions.
The probability of a head is 1/2.
The probability of an odd number is 1/2.
1/2 x 1/2 = 1/4 or 0.25
What occurs when conducting a probability experiment?
An experiment is carried out repeatedly. The total number of times the experiment is conducted and the number of times in which the results are outcomes of interest are recorded. These counts are then used to calculate the experimental probabilities of the outcomes.
What is the most chance of boy baby of intercource in month of the year?
There is no evidence to suggest that the month of intercourse can affect the baby's gender.
How many number combination in the number 1194?
Just one. In combinatorials, the order of the digits in a combination does not make a difference.
Just one. In combinatorials, the order of the digits in a combination does not make a difference.
Just one. In combinatorials, the order of the digits in a combination does not make a difference.
Just one. In combinatorials, the order of the digits in a combination does not make a difference.
probability of pulling out a purple marble = 20/85
probability of NOT pulling out a purple marble = 1 - 20/85 = 65/85 = 13/17
Theoretical probability- tossing 2 fair coins and having 1 land on tails and 1 land on heads?
If it were the 1st one being tails and the 2nd one being heads, then you multiply the probability of each. This is 1/2 times 1/2, which is 1/4; however, since it doesn't specify first and second, that means there is also the favorable outcome of the 1st one being heads and the 2nd one tails, which is another 1/4, for a total of 1/2, or 50%. This makes sense, since all the possible outcomes are H-H, H-T, T-H, and T-T, so having a heads and a tails is 2/4, or 1/2, or 50%.
What are the odds rolling a number less than 3 on a dice.?
-- If you roll one single 6-sided cube . . . probability = 331/3%, odds = 1 in 3, or 2 to 1 against it.
-- If you roll a pair of 6-sided cubes . . . probability = 27/9%, odds = 1 in 36, or 35 to 1 against it.
Examples of relative frequency?
I finished ace combat x on ace but i didn't get the fenrir why?
1.you didn't buy it
2.you didn't get the Falken first
What is the difference between Monte Carlo and stochastic methods of simulation?
monte carlo simulation is used to give solutions of deterministic problems whereas
stochastic simulation is used for stochastic problems.
What is the probability of picking the Jack of Clubs from a pack of cards?
If there is 1 Jack of clubs, and 52 total cards, then the chance of picking the jack on the first selection is 1/52
What percent chance is there of a couple having three daughters in a row?
Each time they have a child, there is a 50% chance it will be female.
Therefore the chance of getting three daughters in a row is 12.5% (0.5 X 0.5 X 0.5 = 0.125).
What is the probability of picking the king queen and jack of hearts from a pack of cards?
The answer depends on how many cards are picked and whether or not the cards are replaced before picking the next one.
If only three cards are picked and they are not replaced, the probability is
3*2*1/(52*51*50) = 1/22100 = 0.000 045 2
If the cards are replace, the probability is
3*2*1/523 = 0.000 042 7